OpenStudy (anonymous):

What is -3/8 ÷ 7/12 A. -7/12 B. -9/14 C. -14/9 D. -32/7 fan and medal. @Queen_Elsa231

OpenStudy (animal_lover36):

= -3/8 ÷ 7/12 = (-3 × 12) / (7 × 8) = -36/56 = -9/14